# Relativistic heavy-ion collisions; classical Yang-Mills theory; glasma

### 09 Jan 2024 v1

@@ 1 Introduction
Relativistic heavy-ion collision experiments in nuclear physics aim to study various states governed by Quantum Chromodynamics (QCD), which describes the strong interaction within the standard model of particle physics. The collision deposits matter at extremely high energy and/or density in the collision region. Subsequently, this matter is believed to evolve into quark-gluon plasma (QGP), where quarks and gluons, normally confined within protons and neutrons in vacuum, are free to move independently. Recently, research has shifted from merely confirming the creation of QGP to exploring its properties through experiments. This includes investigating the effects of magnetic fields B1 ; B2 ; B3 ; B4 ; B5 ; B6 ; B7 ; B8 ; B9 ; B10 ; B11 ; B12 ; B13 ; B14 ; B15 ; B16 ; B17 ; B18 and rotation in the QGP medium omega1 ; omega2 ; omega3 ; omega4 , as well as the observability of novel phenomena related to quantum anomalies, such as the chiral magnetic effect B6 ; B7 and chiral vortical effect CVE1 ; CVE2 ; CVE3 ; CVE4 . We would like to emphasize that heavy-ion collision experiments provide not only the QGP, but also other rich physics landscapes of QCD: including the non-equilibrium state before the formation of QGP and the phase transition from the QGP to the hadron gas.
@@
The pre-equilibrium process between the non-equilibrium state initially produced by the collision and the QGP following relativistic hydrodynamics, remains unclear. Since it is not trivial whether the initial matter with high energy density reaches the QGP before undergoing the QCD phase transition, due to the decrease in energy density caused by the rapid expansion of the created matter, a deeper understanding of the pre-equilibrium process reinforces our confidence that QGP is produced in experiments. Not only that, but it is also vital for explaining the experimental results. For instance, improved modeling of the pre-equilibrium process can provide more accurate initial conditions for the hydrodynamic simulations used in analyses of the experimental data. Additionally, photons emitted from the pre-equilibrium matter may carry information about the non-equilibrium state, as they are less affected by the medium. Of course, the non-equilibrium state of QCD matter itself is also an interesting subject of study.
@@
In the study of non-equilibrium processes in heavy-ion collisions, a key issue is understanding how the created matter eventually behaves as a fluid governed by hydrodynamics. If the system were confined within a box, it might be expected to eventually reach equilibrium over time due to the interactions among the quarks and gluons. However, in reality, the system expands at nearly the speed of light. Therefore, it is not an obvious question whether the created matter can even approach a hydrodynamic state in the first place.
@@
Tracking the real-time evolution of the non-equilibrium state of QCD matter based on first-principle calculation is not practical with our existing computational resources. Therefore, several effective descriptions have been employed to reveal its qualitative characteristics. For example, for the early stage of the collision, the classical Yang-Mills (CYM) field is an excellent description of the system, as explained in the next paragraph.
@@
According to perturbative QCD, in a high-energy collision, a nucleus can be considered as a dense state consisting of soft gluons. Here, ’soft’ implies that the momentum carried is only a small fraction of the hadron’s total momentum. The color glass condensate (CGC) effective theory is a suitable description of such a high-energy nucleus CGC . In this theory, a high-energy nucleus is divided into soft partons (gluons) and hard partons, gluons and quarks with higher momentum than the soft gluons MV1 ; MV2 ; MV3 . Soft gluons are described as the CYM fields emitted from hard partons expressed as classical color charge density. The name ’color glass condensate’ has its origins as follows: ’color’ in QCD is a property analogous to electric charge. ’Glass’ is a disordered system and appears static in natural time. Similarly, hard partons are ’frozen’ in time due to the time dilation effects of special relativity, compared to the lifetime of soft gluons, yet can be rearranged by interactions. ’Condensate’ refers to the dense and coherent soft gluons, which are treated as the classical field, akin to a Bose-Einstein condensate. The theoretical description of the CGC is analogous to the Weizsäcker-Williams approximation in electromagnetic dynamics. As a result of the collision between two CGCs, the initial state of the matter is expected to be a dense matter of soft gluons, a state referred to as ’glasma.’ The glasma is well described by the CYM field, which is determined by solving the CYM equations of motion with two classical color sources, each representing the hard partons within different nuclei glasma .
@@
Under the assumption of the infinite momentum limit, the glasma exhibits exact invariance under the boost transformation along the collision axis and, just after the collision, the color electromagnetic fields of such a boost invariant glasma are aligned along the collision axis 2D1 ; 2D2 ; 2D3 ; 2D4 ; 2D5 ; 2D6 ; 2D7 . However, in real relativistic heavy-ion collisions, there should be fluctuations that break the boost invariance. The CYM equation of motion has instabilities that lead to the rapid growth of these fluctuations Instability1 ; Instability2 ; Instability3 ; Instability4 ; Instability5 ; Instability6 ; Instability7 ; Instability8 ; Instability9 ; Instability10 ; Instability11 ; Instability12 ; Instability13 ; Instability14 ; Instability15 ; Instability16 ; Instability17 ; Instability18 ; Instability19 ; Instability20 ; Instability21 ; Instability22 . Interestingly, these instabilities can drive the pressure isotropization Instability14 ; Instability15 ; Instability16 ; Instability17 ; Instability18 ; Instability19 ; Instability20 ; Instability21 ; Pressure and entropy generation Entropy1 ; Entropy2 ; Entropy3 .
@@
Here, a natural question arises regarding the extent to which the system approaches the hydrodynamic regime during the glasma stage. Many studies on this issue focus on the isotropization of its pressure because sufficient isotropy in the pressure is crucial for the system to obey hydrodynamics. As mentioned above, the pressure isotropy is driven by the growth of fluctuations that break boost invariance, due to instabilities of the CYM theory. However, previous studies have not demonstrated the pressure isotropy consistent with hydrodynamic fluid. For instance, some studies Instability14 ; Instability15 ; Instability16 ; Instability17 ; Instability18 ; Instability19 ; Instability20 ; Instability21 have simulated the semiclassical time evolution of the glasma. These include fluctuations that break boost invariance due to minimal quantum uncertainty. They use the classical statistical approximation (CSA) in a regime with much smaller coupling constants than realistic coupling constants. However, these simulations have not achieved sufficient pressure isotropy. Other study 3DJIMWLK , using the JIMWLK equation -- which is the evolution equation for the separation scale between soft and hard partons in the CGC effective theory -- incorporate the initial condition of the glasma that breaks boost invariance. This initial condition are derived from the momentum distribution of soft gluons in a nucleus with finite momentum. These studies also do not show sufficient pressure isotropy. However, before concluding that the glasma does not exhibit the pressure isotropization under the classical approximation, it remains valuable to consider the effects originating from other physical mechanisms that break boost invariance.
@@
The research is significant as it delves into two critical areas: firstly, it is vital for comprehending how closely the glasma reaches a hydrodynamic fluid within the classical approximation, especially in relation to its impact on the evolution of the degree of the pressure isotropy. Secondly, it provides essential insights for developing more realistic models of the initial stages of the collision.
This article is organized as follows: In Sec. 2, we introduce the 3+1D glasma simulation method that we have recently developed 3Dglasma6 . In Sec. 3, we present the results of both transverse and longitudinal pressure during the evolution of the 3D glasma, comparing pressure isotropization across various longitudinal correlation lengths inside the colliding nuclei. Section 4 is devoted to providing a summary and perspectives on our findings.
@@ 2.1 CGC effective thoery for single relativistic nucleus
In the CGC effective theory, effective for large nuclei at relativistic speeds, we treat ’hard’ partons, gluons and quarks with higher momentum than soft gluons, and soft gluons separately. Hard partons are modeled as classical color charge density $\rho$, while soft gluons are treated as the CYM field $A_{\mu}$ radiated from the classical color current $J_{\mu}$, determined by solving the CYM equations of motion with $J$, $[D_{\mu},F^{\mu\nu}]=J^{\nu}$. A covariant derivative is defined as $D_{\mu}\equiv\partial_{\mu}-igA_{\mu}$, and the field strength is defined by the commutator of $D$, $F_{\mu\nu}\equiv i[D_{\mu},D_{\nu}]$. In this context, only the dominant component of the classical color current in the infinite momentum limit is considered, $J^{\mu}=\delta^{\mu+}\rho/g$, where light-cone coordinates are introduced as $x^{\mp}=(t\mp z)/\sqrt{2}$. The dominance of $+$ component in the classical current is evident because the boost transformation in the positive $z$ direction enhance this component, $J=(J^{+},J^{-},\bm{J}_{\perp})\to(e^{\phi}J^{+},e^{-\phi}J^{-},\bm{J}_{\perp})$, where $\phi$ is a rapidity. Now the continuity equation for the classical color current can be simplified, $[D_{\mu},J^{\mu}]=0\to[D_{+},\rho]=[\partial_{+}-igA_{+},\rho]=0$. Under the gauge condition of $A_{+}=A^{-}=0$, the continuity equation becomes $\partial_{+}\rho=0$, and thus $\rho$ should be static in the light-cone time $x^{+}$,
@@
The soft gauge field, a solution of the equation of motion $[D_{\mu},F^{\mu\nu}]=J^{\nu}$, is a functional of $\rho$ and is known as
@@
Here $\rho_{\rm cov}$ denotes the classical color charge density in the covariant gauge condition, and is related to the color charge density in the $A_{-}=0$ gauge condition through the gauge transformation,
@@
which depends on the separation scale $Y$ between hard partons and soft gluons. When the collision energy is not too high, the classical color charge density consists only of valence quarks. In the ideal situation, where a nucleus is extremely dense and extended infinitely in the transverse direction perpendicular to the collision axis, the weight function for such a color charge density consisting of valence quarks is provided as a Gaussian function MV1 ; MV2 ; MV3 ; Kov , known as the McLerran-Venugopalan (MV) model,
@@
$\langle\rho^{a}(x^{-},\bm{x}_{\perp})\rho^{b}(x^{\prime-},\bm{x}_{\perp}^{\prime})\rangle$
$=\delta^{a,b}\left(g^{2}\mu(x^{-},\bm{x}_{\perp})\right)^{2}\delta(x^{-}-x^{\prime-})\delta^{2}(\bm{x}_{\perp}-\bm{x}_{\perp}^{\prime})\ .$ (8)
As the collision energy increases, $\rho$ should incorporate contributions from an increasing number of particles with large momentum, which can be captured by solving the JIMWLK equation JIMWLK1 ; JIMWLK2 ; JIMWLK3 ; JIMWLK4 ; JIMWLK5 ; JIMWLK6 ; JIMWLK7 ; JIMWLK8 , the evolution equation for the separation scale $Y$, with an initial condition based on the MV model. In the actual application for studying observables in the CGC, we randomly generate an event-by-event color charge density according to the weight function $W_{Y}$ and estimate expectation values of observables as event average values.
Later, we consider the color charge density based on the MV model, which is relevant for the energy of Au-Au collisions at RHIC, with a center-of-mass energy per nucleon pair of 200 GeV. Effects beyond the MV model, which can be taken into account by solving JIMWLK equation, become more important when analyzing larger energy regions at LHC.
@@ 2.2 $3$D glasma in collision of relativistic nuclei
However, the problem becomes much more complicated from the beginning time of the collision. After the collision happens, the CYM field includes the glasma emerging as a result of the collision. Amazingly, the solution representing the glasma is known under the eikonal approximation. This approximation includes three treatments of the classical color charge density or current: firstly, we keep the treatment considering only the dominant components of the classical color current, as in the single nucleus case, which is already assumed in Eq. (9); secondly, we treat the classical color charge density as infinitely thin along the collision axis due to Lorentz contraction, $\rho^{(1/2)}\propto\delta(x^{\mp})$; and thirdly, we assume the classical color charge density is unchanged by the collisions, which means two classical color charge densities are static over the entire time, ±$\partial_{\pm}\rho^{(1/2)}=0$. Under the eikonal approximation, a regular solution of the CYM equation of motion $[D_{\mu},F^{\mu\nu}]=J^{\mu}$ at the infinitely small proper time, $\tau=0+$, is explicitly known as glasma
$A_{i}$ $=\lim_{x^{-}\to\infty}A^{(1)}_{i}+\lim_{x^{-}\to\infty}A^{(2)}_{i}\ ,\ \ A_{\eta}=0\ ,$ (14)
@@
where the Fock-Schwinger gauge condition $A^{\tau}=0$ is imposed, and the electric fields in this gauge are defined as $E^{i}\equiv(\partial_{\tau}A_{i})/\tau$ and $E^{\eta}\equiv\tau\partial_{\tau}A_{\eta}$. Here, Milne coordinates are introduced as $\left(\tau,\eta\right)=\left(\sqrt{2x^{-}x^{+}},\frac{1}{2}\ln{\frac{x^{+}}{x^{-}}}\right)$, which are natural coordinates for describing the matter created in relativistic heavy-ion collisions. The relativistic nucleus exists around $\tau=0$, and the evolution of the created matter expanding along the collision axis at close to the speed of light can be naturally followed in terms of proper time $\tau$, instead of ordinary time $t$.
@@
Some important points about the solution for $\tau=0+$ under the eikonal approximation, shown in Eqs. (14) and (15), should be mentioned. First, this solution represents only the glasma, since the gauge fields representing the soft gluon, the Weizsäcker-Williams field, are located solely at $\tau=0$ in the limit of the infinite thin nucleus. The subsequent evolution of the glasma at a larger $\tau$ can be tracked by solving the CYM equation of motion, using the $\tau=0+$ solution as the initial condition. Second, this solution exhibits exact boost invariance, independent of rapidity, $\partial_{\eta}A=0$ and $\partial_{\eta}E=0$. Third, the color electromagnetic fields in this boost-invariant glasma orient along the collision axis, $E^{\eta},B^{\eta}\neq 0$ and $E^{i},B^{i}=0$. These anisotropic color electromagnetic fields lead to that the longitudinal pressure and the transverse pressure, defined as $P_{\rm\perp}\equiv\tau^{2}T^{\eta\eta}$ and $P_{\perp}\equiv(T^{11}+T^{22})/2$, have the same absolute value but opposite signs, $P_{\rm\perp}=-P_{\rm L}$.
@@
Recently, much attention has been paid to the $3+1$D glasma simulation, which goes beyond the eikonal approximation. In simulation methods proposed in previous papers 3Dglasma1 ; 3Dglasma2 ; 3Dglasma3 ; 3Dglasma4 ; 3Dglasma5 ; 3Dglasma6 , the eikonal approximation is relaxed, allowing the classical color charge density to possess a finite thickness and evolve dynamically according to the continuity equation. Among these methods, we employ the 3+1D glasma simulation method that we have recently developed 3Dglasma6 . The distinctive feature of our $3+1$D glasma simulation method is its use of Milne coordinates to track the evolution of the $3$D glasma, whereas other simulation methods employ Minkowski coordinates. In what follows, we review our $3+1$D glasma simulation method in Milne coordinates.
@@
In the $3+1$D glasma model, the color charge density is no longer assumed to be infinitely thin but has a finite thickness in the direction of the collision axis. Additionally, the color charge density dynamically evolves according to the continuity equation for each nucleus. Since it is not practical to solve a set of CYM equations of motion and continuity equations analytically without any assumptions, we numerically solve these equations in Milne coordinates with the initial collision conditions set for the proper time before the collision occurs. It is important to note that, in this method, we set the center positions of each nucleus in light-cone coordinates not around $x^{\mp}=0$ but around positive finite values, $x^{-}=x^{(1)}_{\rm c}>0$ and $x^{+}=x^{(2)}_{\rm c}>0$. The model parameters, $x^{(1)}_{\rm c}$ and $x^{(2)}_{\rm c}$, can be any value as long as they are sufficiently large. Consequently, the two nuclei do not yet make contact at the proper time of $0+$. To distinguish these Milne coordinates from the usual ones, which are defined so that the centers of mass for each nucleus coincide at $\tau=0$, we introduce new notation, $\left(\tilde{\tau},\tilde{\eta}\right)=\left(\sqrt{2x^{-}x^{+}},\frac{1}{2}\ln{\frac{x^{+}}{x^{-}}}\right)$. Accordingly, the definition of the usual Milne coordinates is now changed to $\left(\tau,\eta\right)=\left(\sqrt{2(x^{-}-x^{(1)}_{\rm c})(x^{+}-x^{(2)}_{\rm c})},\frac{1}{2}\ln{\frac{x^{+}-x^{(2)}_{\rm c}}{x^{-}-x^{(1)}_{\rm c}}}\right)$.
@@
The initial condition of the modified proper time used in the 3+1D glasma simulation, $\tilde{\tau}_{\rm ini}$, is set to be significantly earlier than the moment when the centers of mass for each nucleus coincide. Causality implies that the initial condition of the classical color charge density and CYM field at $\tilde{\tau}=\tilde{\tau}_{\rm ini}$ is given by the sum of the solutions of the evolution equations for each individual nucleus. As shown in Eq. (1), the classical color charge density before the collision is static in the light-cone time. Therefore, the initial classical color charge densities are the same as those in the infinite past,
@@
The way of giving the classical color charge density in the infinite past is explained in the next section. The initial CYM field is given by the sum of the solutions of the CYM equation of motion for each individual nucleus, shown in Eq. (2),

### 29 Feb 2024 v3

@@ 1 Introduction
Relativistic heavy-ion collision experiments in nuclear physics aim to study various states governed by quantum chromodynamics (QCD), which describes the strong interaction within the standard model of particle physics. The collision deposits matter at extremely high energy and/or density in the collision region. Subsequently, this matter is believed to evolve into quark–gluon plasma (QGP), where quarks and gluons, normally confined within protons and neutrons in vacuum, are free to move independently. Recently, research has shifted from merely confirming the creation of QGP to exploring its properties through experiments. This includes investigating the effects of magnetic fields B1 ; B2 ; B3 ; B4 ; B5 ; B6 ; B7 ; B8 ; B9 ; B10 ; B11 ; B12 ; B13 ; B14 ; B15 ; B16 ; B17 ; B18 and rotation in the QGP medium omega1 ; omega2 ; omega3 ; omega4 , as well as the observability of novel phenomena related to quantum anomalies, such as the chiral magnetic effect B6 ; B7 and chiral vortical effect CVE1 ; CVE2 ; CVE3 ; CVE4 . We would like to emphasize that heavy-ion collision experiments provide not only the QGP but also other rich physics landscapes of QCD, including the non-equilibrium state before the formation of QGP and the phase transition from the QGP to the hadron gas.
@@
The pre-equilibrium process between the non-equilibrium state initially produced by the collision and the QGP following relativistic hydrodynamics remains unclear. Since it is not trivial whether the initial matter with high energy density reaches the QGP before undergoing the QCD phase transition, due to the decrease in energy density caused by the rapid expansion of the created matter, a deeper understanding of the pre-equilibrium process reinforces our confidence that QGP is produced in experiments. Not only that, it is also vital for explaining the experimental results. For instance, improved modeling of the pre-equilibrium process can provide more accurate initial conditions for the hydrodynamic simulations used in the analyses of the experimental data. Additionally, photons emitted from the pre-equilibrium matter may carry information about the non-equilibrium state, as they are less affected by the medium. Of course, the non-equilibrium state of QCD matter itself is also an interesting subject of study.
@@
In the study of non-equilibrium processes in heavy-ion collisions, a key issue is understanding how the created matter eventually behaves as a fluid governed by hydrodynamics. If the system were confined within a box, it might be expected to eventually reach equilibrium over time due to the interactions among the quarks and gluons. However, in reality the system expands at nearly the speed of light. Therefore, it is not an obvious question whether or not the created matter can even approach a hydrodynamic state in the first place.
@@
Tracking the real-time evolution of the non-equilibrium state of QCD matter based on first-principle calculation is not practical with our existing computational resources. Therefore, several effective descriptions have been employed to reveal its qualitative characteristics. For example, for the early stage of the collision, the classical Yang–Mills (CYM) field is an excellent description of the system, as explained in the next paragraph.
@@
According to perturbative QCD, in a high-energy collision a nucleus can be considered as a dense state consisting of soft gluons. Here, ‘soft’ implies that the momentum carried is only a small fraction of the hadron’s total momentum. The color glass condensate (CGC) effective theory is a suitable description of such a high-energy nucleus CGC . In this theory, a high-energy nucleus is divided into soft partons (gluons) and hard partons (gluons and quarks with higher momentum than the soft gluons) MV1 ; MV2 ; MV3 . Soft gluons are described as the CYM fields emitted from hard partons expressed as classical color charge density. The name ‘color glass condensate’ has its origins as follows: ‘Color’ in QCD is a property analogous to electric charge. ‘Glass’ is a disordered system and appears static in natural time. Similarly, hard partons are ‘frozen’ in time due to the time dilation effects of special relativity, compared to the lifetime of soft gluons, yet can be rearranged by interactions. ‘Condensate’ refers to the dense and coherent soft gluons, which are treated as the classical field, akin to a Bose–Einstein condensate. The theoretical description of the CGC is analogous to the Weizsäcker–Williams approximation in electromagnetic dynamics. As a result of the collision between two CGCs, the initial state of the matter is expected to be a dense matter of soft gluons, a state referred to as ‘glasma’. The glasma is well described by the CYM field, which is determined by solving the CYM equations of motion with two classical color sources, each representing the hard partons within different nuclei glasma .
@@
Under the assumption of the infinite momentum limit, the glasma exhibits exact invariance under the boost transformation along the collision axis and, just after the collision, the color electromagnetic fields of such a boost invariant glasma are aligned along the collision axis 2D1 ; 2D2 ; 2D3 ; 2D4 ; 2D5 ; 2D6 ; 2D7 ; 2D8 . However, in real relativistic heavy-ion collisions, there should be fluctuations that break the boost invariance. The CYM equation of motion has instabilities that lead to the rapid growth of these fluctuations Instability1 ; Instability2 ; Instability3 ; Instability4 ; Instability5 ; Instability6 ; Instability7 ; Instability8 ; Instability9 ; Instability10 ; Instability11 ; Instability12 ; Instability13 ; Instability14 ; Instability15 ; Instability16 ; Instability17 ; Instability18 ; Instability19 ; Instability20 ; Instability21 ; Instability22 . Interestingly, these instabilities can drive pressure isotropization Instability14 ; Instability15 ; Instability16 ; Instability17 ; Instability18 ; Instability19 ; Instability20 ; Instability21 ; Pressure and entropy generation Entropy1 ; Entropy2 ; Entropy3 .
@@
Here, a natural question arises regarding the extent to which the system approaches the hydrodynamic regime during the glasma stage. Many studies on this issue focus on the isotropization of its pressure because sufficient isotropy in the pressure is crucial for the system to obey hydrodynamics. As mentioned above, the pressure isotropy is driven by the growth of fluctuations that break boost invariance, due to instabilities of the CYM theory. However, previous studies have not demonstrated the pressure isotropy consistent with hydrodynamic fluid. For instance, some studies Instability14 ; Instability15 ; Instability16 ; Instability17 ; Instability18 ; Instability19 ; Instability20 ; Instability21 ; Attractor1 ; Attractor2 have simulated the semiclassical time evolution of the glasma. These include fluctuations that break boost invariance due to minimal quantum uncertainty. They use classical statistical approximation (CSA) in a regime with much smaller coupling constants than realistic coupling constants. However, these simulations have not achieved sufficient pressure isotropy. Another study 3DJIMWLK , using the JIMWLK equation—which is the evolution equation for the separation scale between soft and hard partons in the CGC effective theory—incorporates the initial condition of the glasma that breaks boost invariance. This initial condition is derived from the momentum distribution of soft gluons in a nucleus with finite momentum. These studies also do not show sufficient pressure isotropy. However, before concluding that the glasma does not exhibit the pressure isotropization under the classical approximation, it remains valuable to consider the effects originating from other physical mechanisms that break boost invariance.
@@
The research is significant as it delves into two critical areas: firstly, it is vital for comprehending how closely the glasma reaches a hydrodynamic fluid within the classical approximation, especially in relation to its impact on the evolution of the degree of the pressure isotropy. While pressure isotropy is a reliable indicator for tracking the thermalization of the system, it is crucial to acknowledge that, even when full pressure isotropy is achieved, the classical field is still far from quantum equilibrium. Classical field theory does not reach quantum thermal distribution but exhibits the Rayleigh–Jeans divergence in its equilibrium state. Therefore, it is not a suitable approximation for a quantum system nearing a thermal state, such as a hydrodynamic fluid. Secondly, it provides essential insights for developing more realistic models of the initial stages of the collision.
This article is organized as follows: In Section 2, we introduce the 3+1D glasma simulation method that we have recently developed 3Dglasma6 . In Section 3, we present the results of both transverse and longitudinal pressure during the evolution of the 3D glasma, comparing pressure isotropization across various longitudinal correlation lengths inside the colliding nuclei. Section 4 is devoted to providing a summary and perspectives on our findings.
@@ 2.1 Cgc Effective Thoery for Single Relativistic Nucleus
In the CGC effective theory, effective for large nuclei at relativistic speeds, we treat ‘hard’ partons, gluons and quarks with higher momentum than soft gluons, and soft gluons separately. Hard partons are modeled as classical color charge density $\rho$, while soft gluons are treated as the CYM field $A_{\mu}$ radiated from the classical color current $J_{\mu}$, determined by solving the CYM equations of motion with $J$, $[D_{\mu},F^{\mu\nu}]=J^{\nu}$. A covariant derivative is defined as $D_{\mu}\equiv\partial_{\mu}-igA_{\mu}$, and the field strength is defined by the commutator of $D$, $F_{\mu\nu}\equiv i[D_{\mu},D_{\nu}]$. In this context, only the dominant component of the classical color current in the infinite momentum limit is considered, $J^{\mu}=\delta^{\mu+}\rho/g$, where light-cone coordinates are introduced as $x^{\mp}=(t\mp z)/\sqrt{2}$. The dominance of the $+$ component in the classical current is evident because the boost transformation in the positive $z$ direction enhance this component, $J=(J^{+},J^{-},\bm{J}_{\perp})\to(e^{\phi}J^{+},e^{-\phi}J^{-},\bm{J}_{\perp})$, where $\phi$ is a rapidity. Now the continuity equation for the classical color current can be simplified, $[D_{\mu},J^{\mu}]=0\to[D_{+},\rho]=[\partial_{+}-igA_{+},\rho]=0$. Under the gauge condition of $A_{+}=A^{-}=0$, the continuity equation becomes $\partial_{+}\rho=0$, and thus $\rho$ should be static in the light-cone time $x^{+}$,
@@
The soft gauge field, a solution of the equation of motion, $[D_{\mu},F^{\mu\nu}]=J^{\nu}$, is a functional of $\rho$ and is known as
@@
Here, $\rho_{\rm cov}$ denotes the classical color charge density in the covariant gauge condition and is related to the color charge density in the $A_{-}=0$ gauge condition through the gauge transformation,
@@
which depends on the separation scale $Y$ between hard partons and soft gluons. When the collision energy is not too high, the classical color charge density consists only of valence quarks. In the ideal situation, where a nucleus is extremely dense and extended infinitely in the transverse direction perpendicular to the collision axis, the weight function for such a color charge density consisting of valence quarks is provided as a Gaussian function MV1 ; MV2 ; MV3 ; Kov , known as the McLerran–Venugopalan (MV) model,
@@
$\begin{split}&\langle\rho^{a}(x^{-},\bm{x}_{\perp})\rho^{b}(x^{\prime-},\bm{x}_{\perp}^{\prime})\rangle\\ &=\delta^{a,b}\left(g^{2}\mu(x^{-},\bm{x}_{\perp})\right)^{2}\delta(x^{-}-x^{\prime-})\delta^{2}(\bm{x}_{\perp}-\bm{x}_{\perp}^{\prime})\ .\end{split}$ (8)
As the collision energy increases, $\rho$ should incorporate contributions from an increasing number of particles with large momentum, which can be captured by solving the JIMWLK equation JIMWLK1 ; JIMWLK2 ; JIMWLK3 ; JIMWLK4 ; JIMWLK5 ; JIMWLK6 ; JIMWLK7 ; JIMWLK8 , the evolution equation for the separation scale, $Y$, with an initial condition based on the MV model. In the actual application for studying observables in the CGC, we randomly generate an event-by-event color charge density according to the weight function, $W_{Y}$, and estimate the expectation values of observables as event average values.
Later, we consider the color charge density based on the MV model, which is relevant for the energy of Au–Au collisions at RHIC, with a center-of-mass energy per nucleon pair of 200 GeV. Effects beyond the MV model, which can be taken into account by solving the JIMWLK equation, become more important when analyzing larger energy regions at LHC.
@@ 2.2 $3$D Glasma in Collision of Relativistic Nuclei
However, the problem becomes much more complicated from the beginning time of the collision. After the collision happens, the CYM field includes the glasma emerging as a result of the collision. Amazingly, the solution representing the glasma is known under the eikonal approximation. This approximation includes three treatments of the classical color charge density or current: firstly, we keep the treatment considering only the dominant components of the classical color current, as in the single nucleus case, which is already assumed in Equation (9); secondly, we treat the classical color charge density as infinitely thin along the collision axis due to Lorentz contraction, $\rho^{(1/2)}\propto\delta(x^{\mp})$; and thirdly, we assume the classical color charge density is unchanged by the collisions, which means two classical color charge densities are static over the entire time, ±$\partial_{\pm}\rho^{(1/2)}=0$. Under the eikonal approximation, a regular solution of the CYM equation of motion, $[D_{\mu},F^{\mu\nu}]=J^{\mu}$, at the infinitely small proper time, $\tau=0+$, is explicitly known as glasma
$A_{i}$ $=\lim_{x^{-}\to\infty}A^{(1)}_{i}+\lim_{x^{+}\to\infty}A^{(2)}_{i}\ ,\ \ A_{\eta}=0\ ,$ (14)
@@
where the Fock–Schwinger gauge condition $A^{\tau}=0$ is imposed, and the electric fields in this gauge are defined as ${E^{i}\equiv\tau\partial_{\tau}A_{i}}$ and ${E^{\eta}\equiv\partial_{\tau}A_{\eta}/\tau}$. Here, Milne coordinates are introduced as $\left(\tau,\eta\right)=\left(\sqrt{2x^{-}x^{+}},\frac{1}{2}\ln{\frac{x^{+}}{x^{-}}}\right)$, which are natural coordinates for describing the matter created in relativistic heavy-ion collisions. The relativistic nucleus exists around $\tau=0$, and the evolution of the created matter expanding along the collision axis at close to the speed of light can be naturally followed in terms of proper time, $\tau$, instead of ordinary time, $t$.
@@
Some important points about the solution for $\tau=0+$ under the eikonal approximation, shown in Equations (14) and (15), should be mentioned. First, this solution represents only the glasma, since the gauge fields representing the soft gluon, the Weizsäcker–Williams field, are located solely at $\tau=0$ in the limit of the infinite thin nucleus. The subsequent evolution of the glasma at a larger $\tau$ can be tracked by solving the CYM equation of motion, using the $\tau=0+$ solution as the initial condition. Second, this solution exhibits exact boost invariance, independent of rapidity, $\partial_{\eta}A=0$ and $\partial_{\eta}E=0$. Third, the color electromagnetic fields in this boost-invariant glasma orient along the collision axis, $E^{\eta},B^{\eta}\neq 0$ and $E^{i},B^{i}=0$. These anisotropic color electromagnetic fields lead to the longitudinal pressure and the transverse pressure, defined as $P_{\rm\perp}\equiv\tau^{2}T^{\eta\eta}$ and $P_{\perp}\equiv(T^{11}+T^{22})/2$, having the same absolute value but opposite signs, $P_{\rm\perp}=-P_{\rm L}$.
@@
Recently, much attention has been paid to the 3+1D glasma simulation, which goes beyond the eikonal approximation. In the simulation methods proposed in previous papers 3Dglasma1 ; 3Dglasma2 ; 3Dglasma3 ; 3Dglasma4 ; 3Dglasma5 ; 3Dglasma6 , the eikonal approximation is relaxed, allowing the classical color charge density to possess a finite thickness and evolve dynamically according to the continuity equation. Among these methods, we employ the 3+1D glasma simulation method that we have recently developed 3Dglasma6 . The distinctive feature of our 3+1D glasma simulation method is its use of Milne coordinates to track the evolution of the $3$D glasma, whereas other simulation methods employ Minkowski coordinates. In what follows, we review our 3+1D glasma simulation method in Milne coordinates.
@@
In the 3+1D glasma model, the color charge density is no longer assumed to be infinitely thin but has a finite thickness in the direction of the collision axis. Additionally, the color charge density dynamically evolves according to the continuity equation for each nucleus. Since it is not practical to solve a set of CYM equations of motion and continuity equations analytically without any assumptions, we numerically solve these equations in Milne coordinates with the initial collision conditions set for the proper time before the collision occurs. It is important to note that, in this method, we set the center positions of each nucleus in light-cone coordinates not around $x^{\mp}=0$ but around positive finite values, $x^{-}=x^{(1)}_{\rm c}>0$ and $x^{+}=x^{(2)}_{\rm c}>0$. The model parameters $x^{(1)}_{\rm c}$ and $x^{(2)}_{\rm c}$ can be any value as long as they are sufficiently large. Consequently, the two nuclei do not yet make contact at the proper time of $0+$. To distinguish these Milne coordinates from the usual ones, which are defined so that the centers of mass for each nucleus coincide at $\tau=0$, we introduce new notation, $\left(\tilde{\tau},\tilde{\eta}\right)=\left(\sqrt{2x^{-}x^{+}},\frac{1}{2}\ln{\frac{x^{+}}{x^{-}}}\right)$. Accordingly, the definition of the usual Milne coordinates is now changed to $\left(\tau,\eta\right)=\left(\sqrt{2(x^{-}-x^{(1)}_{\rm c})(x^{+}-x^{(2)}_{\rm c})},\frac{1}{2}\ln{\frac{x^{+}-x^{(2)}_{\rm c}}{x^{-}-x^{(1)}_{\rm c}}}\right)$.
@@
The initial condition of the modified proper time used in the 3+1D glasma simulation, $\tilde{\tau}_{\rm ini}$, is set to be significantly earlier than the moment when the centers of mass for each nucleus coincide. Causality implies that the initial condition of the classical color charge density and CYM field at $\tilde{\tau}=\tilde{\tau}_{\rm ini}$ is given by the sum of the solutions of the evolution equations for each individual nucleus. As shown in Equation (1), the classical color charge density before the collision is static in the light-cone time. Therefore, the initial classical color charge densities are the same as those in the infinite past,
@@
The way of giving the classical color charge density in the infinite past is explained in the next section. The initial CYM field is given by the sum of the solutions of the CYM equation of motion for each individual nucleus, as shown in Equation (2),